In an experiment, a glass tube is submerged into a larger graduated cylinder full of water. A tuning fork is vibrating above the tube at a frequency of 659 Hz. If the 1st resonant frequency is heard at 14.2 cm, what is the speed of sound? Round your answer to the nearest meter.
1
Expert's answer
2020-06-22T11:13:03-0400

The resonant frequency "f" , the lenght of the tube "L" and the speed of the sound "v" are connected with the following expression (see https://en.wikipedia.org/wiki/Acoustic_resonance#Open_at_both_ends):


"f = \dfrac{nv}{2L}"

Expressing "v" from this equation and setting "n=1" (for the 1st resonant frequency), obtain:


"v = 2fL"

Substituting numerical values, get:


"v = 2\cdot 659Hz\cdot 0.142m = 187.156 m/s\approx187m/s"

Answer. v = 187 m/s.